OpenStudy (anonymous):

for what value of k are the roots of x^2+kx+600=0 consecutive integers

OpenStudy (loser66):

We have the property of roots of quadratic \(x_1+x_2=\dfrac{-b}{a}\) and \(x_1*x_2=\dfrac{c}{a}\) just replace \(x_2= x_1+1\) and solve for k